# Global well-posedness for small data in a 3D temperature-velocity model with Dirichlet boundary noise

### v2

@@ 1. Introduction
The parameter $\varepsilon$ influences the probability of global existence up to time $T$. Indeed, the three-dimensional (3D) Navier-Stokes equations are only known to be globally well‑posed under smallness conditions on initial data and forcing (see, e.g., [8, Chapter 9]). Since stochastic boundary forcing enters the velocity equation through buoyancy coupling, we impose a small‑noise regime in order to retain a nonzero probability of well‑posedness on $[0,T]$. However, even for arbitrarily small $\varepsilon>0$, the boundary noise may assume large values with positive probability. Consequently, we construct the solution $(u^{\varepsilon},\theta^{\varepsilon})$ up to a random stopping time $\tau^{\varepsilon}\leq T$, defined to interrupt the evolution if the stochastic forcing becomes too large.
@@
where $s\in[0,1/2)$ is the Sobolev regularity of the initial temperature $\theta_{0}\in W^{s,6/5}(\mathcal{D})$.
@@
Furthermore, we establish the high-probability estimate $\mathbb{P}(\tau^{\varepsilon}=T)\geq 1-C\varepsilon$. As $s\uparrow 1/2$, the constraint from the initial datum disappears and one may choose $\delta_{u}=\delta_{\theta}$, recovering a symmetric regularity regime.
It is worth underlining two aspects of the boundary noise: its mathematical difficulties, and its physical meaning. Firstly, Da Prato and Zabczyk were the first to rigorously demonstrate in [9] that solutions to the heat equation subject to Dirichlet boundary noise exhibit significantly reduced regularity compared to solutions driven by interior stochastic forcing, even in one spatial dimension. Specifically, due to the presence of boundary noise, the best regularity that one can hope for, even in the linear case, is continuity in time with values in a negative-order Bessel potential space, namely $C_{t}(H_{x}^{-\frac{1}{2}-\delta_{\theta}})$ for any $\delta_{\theta}>0$. This limitation reflects the difficulties of noisy Dirichlet boundary conditions.
@@
In parallel, the velocity field $u^{\varepsilon}$ enjoys, in this context, the maximal regularity of the Stokes operator in the same low-regularity framework introduced above. This constitutes the optimal regularity according to the maximal regularity theory for the Stokes operator on $H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D})$, with suitable divergence-free and boundary conditions on the spaces. Our maximal regularity approach follows the framework developed by Prüss and Wilke in [25, Section 5].
Secondly, from a physical point of view, stochastic boundary forcing encapsulates the influence of unresolved, rapid fluctuations, such as boundary-layer instabilities, or small-scale convection, that cannot be parametrized deterministically at the scale of our model. This philosophy traces back to Hasselmann’s stochastic climate paradigm [18], in which slow, large-scale dynamics are driven by fast, random perturbations. The boundary-noise framework thus provides a mathematically tractable way to “close” the system. In this spirit, we plan to investigate in a future work in which sense the system (1.1) arises as a limit of a multiscale fast-slow model.
@@ 2.1. Notations and functional setting
$W(t,x)=\sum_{k\geq 0}\lambda_{k}e_{k}(x)\beta_{k}(t),\quad t\geq 0,\ x\in\partial\mathcal{D},$
@@
where $H^{2\alpha}_{00}(\mathcal{D})$ consists of all $u\in H^{\frac{1}{2}}(\mathcal{D})$ such that
@@
In the Hilbert space $X_{0}=L^{2}_{\sigma}(\mathcal{D})$, the Stokes operator $A$ is a non-negative, self-adjoint operator with compact inverse; see, for instance, [27, Section 2]. In particular, $A$ is sectorial of angle 0 and A-A generates a bounded analytic $C_{0}$-semigroup on $X_{0}$; see, for example, [12, 19]. Moreover, since $A$ is a non-negative self-adjoint operator on a Hilbert space, it admits a bounded $\mathcal{H}^{\infty}$-functional calculus of angle 0; see [17, Chapters 2 and 7]. In the specific case of the Stokes operator, the domains of the fractional powers $A^{\alpha}$ in $L^{2}_{\sigma}(\mathcal{D})$ have been identified in [15] and are given by
@@ 2.2. Description of our approach
We conclude this section by describing the solution of the linear temperature problem associated to (2.2), i.e. by recalling the following result concerning the mild solution of (2.4), which follows from [9] and the factorisation trick for the stochastic convolution [11, Section 5.3.1].
@@ 2.3. Main results
We are now in a position to state the main results of our work. The first one concerns the well-posedness of the system (2.5).
@@
Let $s\in[0,1/2)$ and $p>2$. Then there exist $\overline{\delta}_{\theta}=\overline{\delta}_{\theta}(s,p)>0$ and $\overline{\delta}_{u}=\overline{\delta}_{u}(s,p)>0$ such that for every $0<\delta_{\theta}\leq\overline{\delta}_{\theta}$ and $0<\delta_{u}<\overline{\delta}_{u}$, the following holds. Let
$\alpha_{\theta}=\frac{1}{4}+\frac{\delta_{\theta}}{2},\quad\beta_{\theta}=\frac{1}{4}-\frac{\delta_{\theta}}{4},$
and denote by $Z^{\varepsilon}_{t}=-\sqrt{\varepsilon}\int_{0}^{t}\Delta e^{(t-r)\Delta}D\,dW_{r}$ the continuous version in $H^{-2\alpha_{\theta}}(\mathcal{D})$ considered in Proposition 2.1. Then, for every divergence–free velocity field
@@
$\zeta^{\varepsilon}\in C(0,T;W^{s,6/5}(\mathcal{D}))\quad\mathbb{P}-\text{a.s.}$
@@
We now turn to the analysis of the 3D Navier-Stokes equations (2.3), assuming that $f^{\theta}$ is a prescribed forcing term. Global well-posedness results for the 3D Navier–Stokes equations are well known under smallness assumptions on the initial datum and the forcing; see for instance [8, Chapter 9]. In the classical energy setting one typically assumes $f^{\theta}\in L^{2}(0,T;L^{2}(\mathcal{D}))$ and $u_{0}\in H^{1}(\mathcal{D})$.
@@
The triple $(u^{\varepsilon},\theta^{\varepsilon},\tau^{\varepsilon})$ is a solution of (1.1) on $[0,T]$ if $\tau^{\varepsilon}\colon\Omega\to[0,T]$ is a stopping time and $(u^{\varepsilon},\theta^{\varepsilon})$ is a stochastic process with trajectories in
@@
Let $T>0$ and $s\in[0,\tfrac{1}{2})$. Fix $p>4.$ Choose a parameter $\gamma>\frac{1}{4}$ sufficiently close to $\frac{1}{4}$ such that
$a(s,\gamma):=\frac{1}{2}+\gamma+\frac{s}{2}<1,\qquad\lambda_{\max}:=p(1-a(s,\gamma))>\lambda_{\min}:=\frac{1}{2}-s.$
(Such a $\gamma$ always exists since $p>4$ implies p(1a(s,1/4))>1/2sp(1-a(s,1/4))>1/2-s). Choose
$\lambda\in(\lambda_{\min},\,\lambda_{\max})\cap(0,1),$
and define the thresholds
$\overline{\delta}_{\theta}:=\min\{\lambda,\ 2\gamma-\tfrac{1}{2}\},\quad\overline{\delta}_{u}:=\min\{\lambda_{\max}-\lambda,1-\lambda\}.$
Let parameters $\delta_{\theta},\delta_{u}$ satisfy
$0<\delta_{\theta}\leq\overline{\delta}_{\theta},\qquad 0<\delta_{u}<\overline{\delta}_{u},$
@@
$\delta_{u}\geq\max\left\{\delta_{\theta},\ \tfrac{1}{2}-s\right\},\quad\delta_{u}<1-\frac{2}{p}.$
Set
$\alpha_{\theta}:=\tfrac{1}{4}+\tfrac{\delta_{\theta}}{2},\qquad\beta_{\theta}:=\tfrac{1}{4}-\tfrac{\delta_{\theta}}{4},$
@@
Remark 2.9 (Consistency of parameters).
Let us briefly verify that the set of admissible parameters is non-empty under the condition $p>4$.
  1. (1)
    Existence of $\lambda$. As $\gamma\downarrow\frac{1}{4}$, we have $a(s,\gamma)\to\frac{3}{4}+\frac{s}{2}$. Consequently, $\lambda_{\max}\to p(\frac{1}{4}-\frac{s}{2})$. The condition $\lambda_{\max}>\lambda_{\min}=\frac{1}{2}-s$ is equivalent to $p>\frac{4(1/2-s)}{1-2s}=2$. Since $p>4$, the interval for $\lambda$ is non-empty for $\gamma$ sufficiently close to $1/4$.
  2. (2)
    Existence of $\delta_{u}$. We need to ensure that the lower bound for $\delta_{u}$ is strictly smaller than the upper bound. Taking $\delta_{\theta}$ sufficiently small, the lower bound is $\frac{1}{2}-s$. On the other hand, the upper bound is $\min\{\lambda_{\max}-\lambda,1-\lambda,1-\frac{2}{p}\}$. We thus need to check that there exists a $\lambda$ such that
    $\frac{1}{2}-s<\min\left\{1-\frac{2}{p},\ \lambda_{\max}-\lambda,\ 1-\lambda\right\}.$
    • First, the condition $\frac{1}{2}-s<1-\frac{2}{p}$ is equivalent to $p>\frac{4}{1+2s}$. Since $s\in[0,1/2)$, the term $\frac{4}{1+2s}$ is at most $4$. Thus, the hypothesis $p>4$ guarantees this condition holds.
    • Second, we check that $\frac{1}{2}-s<\lambda_{\max}-\lambda$. This is equivalent to finding $\lambda$ such that $\lambda<\lambda_{\max}-(\frac{1}{2}-s)$. Since we can choose $\lambda$ arbitrarily close to its lower bound $\lambda_{\min}=\frac{1}{2}-s$, such a $\lambda$ exists if $\lambda_{\min}<\lambda_{\max}-\lambda_{\min}$, i.e., $2\lambda_{\min}<\lambda_{\max}$. Since $\lambda_{\max}=p\bigl(1-a(s,\gamma)\bigr)$ and $a(s,\gamma)\to\frac{3}{4}+\frac{s}{2}$ as $\gamma\downarrow\frac{1}{4}$, we have $\lim\limits_{\gamma\downarrow\frac{1}{4}}\lambda_{\max}=\frac{p}{4}(1-2s);$ note that $\frac{p}{4}(1-2s)>2\lambda_{\min}=1-2s$, thanks to the hypothesis $p>4$. Thus the original condition is verified.
    • Finally, we check that $\frac{1}{2}-s<1-\lambda$. This is equivalent to finding $\lambda$ such that $\lambda<\frac{1}{2}+s$. Since $\lambda$ can be chosen close to $\lambda_{\min}=\frac{1}{2}-s$, i.e. $\lambda=\lambda_{\min}+\varepsilon_{\lambda},$ with $\varepsilon_{\lambda}>0$. Thus the condition can be rewritten as $\frac{1}{2}-s<\frac{1}{2}+s+\varepsilon_{\lambda}$, which holds for any $s\geq 0$.
@@ 3.1. Stochastic linear problem with non-homogeneous Dirichlet boundary conditions
We start this section by proving the first part of Proposition 2.1, in particular the well-posedness and the regularity for the mild solution $Z_{t}^{\varepsilon}$ of the linear problem (2.4) subject to Dirichlet noise boundary conditions. Note that this result is well-known in the literature, see for instance [9, Proposition 3.1].
@@
then the thesis follows from the Kolmogorov test for Gaussian processes, see [11, Proposition 3.16]. Consider
@@
For the estimate of the second integral $I_{2}$, we use the fact that for any $\gamma\in(0,\alpha_{\theta}+\beta_{\theta})$ there exists $C>0$ such that for any $0\leq r\leq t_{1}$, it holds
@@
(3.5) $\|(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}(e^{(t_{2}-r)\Delta}-e^{(t_{1}-r)\Delta})\|_{\mathcal{L}(L^{2}(\mathcal{D}))}\leq C(t_{2}-t_{1})^{\gamma}(t_{1}-r)^{-(1-\alpha_{\theta}-\beta_{\theta}+\gamma)},$
where $\mathcal{L}(L^{2})$ is the space of bounded, linear operators from $L^{2}(\mathcal{D})$ to itself. Indeed, by the semigroup property
$e^{(t_{2}-r)\Delta}-e^{(t_{1}-r)\Delta}=e^{(t_{1}-r)\Delta}(e^{(t_{2}-t_{1})\Delta}-I).$
Further, by the commuting property of the fractional powers with the heat semigroup
$(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}(e^{(t_{2}-r)\Delta}-e^{(t_{1}-r)\Delta})=(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}+\gamma}e^{(t_{1}-r)\Delta}(-\Delta)^{-\gamma}(e^{(t_{2}-t_{1})\Delta}-I).$
Considering the norms on the previous identity, it yields
(3.6) $\|(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}(e^{(t_{2}-r)\Delta}-e^{(t_{1}-r)\Delta})\|_{\mathcal{L}(L^{2})}\leq\|(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}+\gamma}e^{(t_{1}-r)\Delta}\|_{\mathcal{L}(L^{2})}\|(-\Delta)^{-\gamma}(e^{(t_{2}-t_{1})\Delta}-I)\|_{\mathcal{L}(L^{2})}.$
The first term on the right-hand-side can be estimated by the analyticity of the semigroup as
(3.7) $\|(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}+\gamma}e^{(t_{1}-r)\Delta}\|_{\mathcal{L}(L^{2})}\leq C(t_{1}-r)^{-1+\alpha_{\theta}+\beta_{\theta}-\gamma}.$
For the second term on the right-hand side, again thanks to the analyticity of the semigroup, it holds
$(e^{(t_{2}-t_{1})\Delta}-I)g=\Delta\int_{0}^{t_{2}-t_{1}}e^{\sigma\Delta}g\,d\sigma$
for any $g\in L^{2}(\mathcal{D})$, see [22, Proposition 2.1.4]. Applying $(-\Delta)^{-\gamma}$, using $1-\gamma<1$, the fact that $(-\Delta)^{1-\gamma}$ is a closed operator and $\sigma\mapsto(-\Delta)^{1-\gamma}e^{\sigma\Delta}$ is integrable for any $g\in L^{2}(\mathcal{D})$, we obtain, thanks to [21, Proposition A.5],
$(-\Delta)^{-\gamma}(e^{(t_{2}-t_{1})\Delta}-I)g=-\int_{0}^{t_{2}-t_{1}}(-\Delta)^{1-\gamma}e^{\sigma\Delta}g\,d\sigma.$
Again by the estimate for the fractional powers, we have
$\begin{split}\|(-\Delta)^{-\gamma}(e^{(t_{2}-t_{1})\Delta}-I)g\|_{2}&\leq\int_{0}^{t_{2}-t_{1}}\|(-\Delta)^{1-\gamma}e^{\sigma\Delta}g\|_{2}\,d\sigma\\ &\leq C\int_{0}^{t_{2}-t_{1}}\sigma^{-1+\gamma}\|g\|_{2}\,d\sigma\\ &=\frac{C}{\gamma}(t_{2}-t_{1})^{\gamma}\|g\|_{2}.\end{split}$
Since the previous estimate holds for any $g\in L^{2}(\mathcal{D})$, we have checked
(3.8) $\|(-\Delta)^{-\gamma}(e^{(t_{2}-t_{1})\Delta}-I)\|_{\mathcal{L}(L^{2})}\leq\frac{C}{\gamma}(t_{2}-t_{1})^{\gamma}$
for any $\gamma\in(0,\alpha_{\theta}+\beta_{\theta}).$ Substituting (3.7) and (3.8) into (3.6) gives (3.5).
Thus, to estimate $I_{2}$, using the independence of $(\beta_{k})_{k}$, Itô-isometry and (3.5), we obtain, for any $\gamma\in(0,\alpha_{\theta}+\beta_{\theta})$,
(3.9) $\begin{split}\mathbb{E}\left[\|I_{2}\|_{2}^{2}\right]&\leq\sum_{k}\lambda_{k}^{2}\int_{0}^{t_{1}}\|(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}\left[e^{(t_{2}-r)\Delta}-e^{(t_{1}-r)\Delta}\right](-\Delta)^{\beta_{\theta}}De_{k}\|_{2}^{2}\,dr\\ &\leq C\sum_{k}\lambda_{k}^{2}\|(-\Delta)^{\beta_{\theta}}De_{k}\|_{2}^{2}\,(t_{2}-t_{1})^{2\gamma}\int_{0}^{t_{1}}(t_{1}-r)^{-2(1-\alpha_{\theta}-\beta_{\theta}+\gamma)}\,dr.\end{split}$
Note that the integral on the right-hand side is finite if $1-\alpha_{\theta}-\beta_{\theta}+\gamma<\frac{1}{2},$ i.e. $\gamma<\alpha_{\theta}+\beta_{\theta}-\frac{1}{2}=\frac{\delta_{\theta}}{4}$. Thus, choosing $\gamma\in(0,\frac{\delta_{\theta}}{4})$, we arrive to
@@
(3.10) $\mathbb{E}\left[\|I_{2}\|_{2}^{2}\right]\leq C\sum_{k}\lambda_{k}^{2}\|(-\Delta)^{\beta_{\theta}}De_{k}\|_{2}^{2}(t_{2}-t_{1})^{2\gamma},$
@@
By the factorisation method, see [11, Section 5.3.1], since $\int_{0}^{T}t^{-2\rho}\|e^{t\Delta}\Phi\|_{L^{0}_{2}}^{2}dt<\infty$, it holds

### v3

@@ 1. Introduction
The parameter $\varepsilon$ influences the probability of global existence up to time $T$. Indeed, the three-dimensional (3D) Navier-Stokes equations are only known to be globally well‑posed under smallness conditions on initial data and forcing (see, e.g., [9, Chapter 9]). Since stochastic boundary forcing enters the velocity equation through buoyancy coupling, we impose a small‑noise regime in order to retain a nonzero probability of well‑posedness on $[0,T]$. However, even for arbitrarily small $\varepsilon>0$, the boundary noise may assume large values with positive probability. Consequently, we construct the solution $(u^{\varepsilon},\theta^{\varepsilon})$ up to a random stopping time $\tau^{\varepsilon}\leq T$, defined to interrupt the evolution if the stochastic forcing becomes too large.
@@
where $s\in(0,1/2)$ is the Sobolev regularity of the initial temperature $\theta_{0}\in W^{s,6/5}(\mathcal{D})$.
@@
Furthermore, we establish the high-probability estimate $\mathbb{P}(\tau^{\varepsilon}=T)\geq 1-C\varepsilon$.
It is worth underlining two aspects of the boundary noise: its mathematical difficulties, and its physical meaning. Firstly, Da Prato and Zabczyk were the first to rigorously demonstrate in [10] that solutions to the heat equation subject to Dirichlet boundary noise exhibit significantly reduced regularity compared to solutions driven by interior stochastic forcing, even in one spatial dimension. Specifically, due to the presence of boundary noise, the best regularity that one can hope for, even in the linear case, is continuity in time with values in a negative-order Bessel potential space, namely $C(0,T;H^{-\frac{1}{2}-\delta_{\theta}}(\mathcal{D}))$ for any $\delta_{\theta}>0$ and any final time $T>0$. This limitation reflects the difficulties of noisy Dirichlet boundary conditions.
@@
In parallel, the velocity field $u^{\varepsilon}$ enjoys, in this context, the maximal regularity of the Stokes operator in the same low-regularity framework introduced above. This constitutes the optimal regularity according to the maximal regularity theory for the Stokes operator on $H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D})$, with suitable divergence-free and boundary conditions on the spaces. Our maximal regularity approach follows the framework developed by Prüss and Wilke in [27, Section 5].
Secondly, from a physical point of view, stochastic boundary forcing represents the effect of unresolved fast fluctuations, such as boundary-layer instabilities or small-scale convection, that cannot be described deterministically at the scale of our model. Related approaches that model boundary-layer effects through non-standard boundary mechanisms can be found, for instance, in [7]. In climate science, this philosophy traces back to Hasselmann’s stochastic climate paradigm [19], in which slow, large-scale dynamics are driven by fast, random perturbations. The boundary-noise framework thus provides a mathematically tractable way to “close” the system. In this spirit, we plan to investigate in a future work in which sense the system (1.1) arises as a limit of a multiscale fast-slow model.
@@ 2.1. Notations and functional setting
$W(t,x)=\sum_{k\geq 0}\lambda_{k}e_{k}(x)\beta_{k}(t),\quad t\geq 0,\ x\in\partial\mathcal{D}$
@@
where $H^{\frac{1}{2}}_{00}(\mathcal{D})$ consists of all $u\in H^{\frac{1}{2}}(\mathcal{D})$ such that
@@
In the Hilbert space $X_{0}=L^{2}_{\sigma}(\mathcal{D})$, the Stokes operator $A$ is a non-negative, self-adjoint operator with compact inverse; see, for instance, [29, Section 2]. In particular, $A$ is sectorial of angle 0 and A-A generates a bounded analytic $C_{0}$-semigroup on $X_{0}$; see, for example, [13, 20]. Moreover, since $A$ is a non-negative self-adjoint operator on a Hilbert space, it admits a bounded $\mathcal{H}^{\infty}$-functional calculus of angle 0; see [18, Chapters 2 and 7]. In the specific case of the Stokes operator, the domains of the fractional powers $A^{\alpha}$ in $L^{2}_{\sigma}(\mathcal{D})$ have been identified in [16] and are given by
@@ 2.2. Description of our approach
We conclude this section by describing the solution of the linear temperature problem associated to (2.2), i.e. by recalling the following result concerning the mild solution of (2.4), which follows from [10] and the factorisation trick for the stochastic convolution [12, Section 5.3.1].
@@ 2.3. Main results
We are now in a position to state the main results of our work. For this purpose, we begin by introducing some notation.
Let $p>4$ and $s\in(0,\tfrac{1}{2})$. Choose
(2.8) $\gamma:=\frac{1}{4}+\varepsilon_{\gamma},\qquad 0<\varepsilon_{\gamma}<\min\left\{\frac{1-2s}{4},\ \frac{(1-2s)(p-4)}{4p}\right\},$
and set $a(s,\gamma):=\frac{1}{2}+\gamma+\frac{s}{2}$ and $\lambda_{\max}:=p(1-a(s,\gamma))$. Then
$a(s,\gamma)=\frac{3}{4}+\frac{s}{2}+\varepsilon_{\gamma}<1\quad\text{(since }\varepsilon_{\gamma}<\tfrac{1-2s}{4}\text{)},$
and, using $1-a(s,\gamma)=\frac{1-2s}{4}-\varepsilon_{\gamma}$,
(2.9) $\lambda_{\max}-\left(\frac{1}{2}-s\right)=p\left(\frac{1-2s}{4}-\varepsilon_{\gamma}\right)-\frac{1-2s}{2}=\frac{(1-2s)(p-2)}{4}-p\varepsilon_{\gamma}>0,$
so $\lambda_{\max}>\frac{1}{2}-s$. Moreover,
(2.10) $\lambda_{\max}-(1-2s)=p\left(\frac{1-2s}{4}-\varepsilon_{\gamma}\right)-(1-2s)=\frac{(1-2s)(p-4)}{4}-p\varepsilon_{\gamma}>0,$
by (2.8). Hence the interval $(0,\min\{2s,\ \lambda_{\max}-(1-2s)\})$ is non-empty, and we may choose
(2.11) $\lambda:=\left(\frac{1}{2}-s\right)+\varepsilon_{\lambda},\qquad 0<\varepsilon_{\lambda}<\min\left\{2s,\lambda_{\max}-(1-2s)\right\}.$
In particular, $\lambda<\frac{1}{2}+s<1$ (since $\varepsilon_{\lambda}<2s$). Finally, define the thresholds
(2.12) $\overline{\delta}_{\theta}:=\min\left\{\lambda,2\gamma-\frac{1}{2}\right\},\qquad\overline{\delta}_{u}:=\min\left\{\lambda_{\max}-\lambda,1-\lambda\right\}.$
The first main result concerns the well-posedness of the system (2.5)
@@
Let $T>0,\,s\in(0,\tfrac{1}{2})$ and $p>4$, and choose $\gamma,\lambda$ as in (2.8) and (2.11). Let $\overline{\delta}_{\theta}(s,p)$ and $\overline{\delta}_{u}(s,p)$ be given by (2.12). Then for every
$0<\delta_{\theta}<\overline{\delta}_{\theta}(s,p),\qquad 0<\delta_{u}<\overline{\delta}_{u}(s,p),$
the following holds. Set
$\alpha_{\theta}=\frac{1}{4}+\frac{\delta_{\theta}}{2},\qquad\beta_{\theta}=\frac{1}{4}-\frac{\delta_{\theta}}{4},$
and let $Z_{t}^{\varepsilon}=-\sqrt{\varepsilon}\int_{0}^{t}\Delta e^{(t-r)\Delta}D\,dW_{r}$ be the continuous version in $H^{-2\alpha_{\theta}}(\mathcal{D})$ from Proposition 2.1. Then, for every divergence-free velocity field
@@
$\zeta^{\varepsilon}\in C(0,T;W^{s,6/5}(\mathcal{D}))\qquad\mathbb{P}\text{-a.s.}$
@@
We now turn to the analysis of the 3D Navier-Stokes equations (2.3), assuming that $f^{\theta}$ is a prescribed forcing term. Global well-posedness results for the 3D Navier–Stokes equations are well known under smallness assumptions on the initial datum and the forcing; see for instance [9, Chapter 9]. In the classical energy setting one typically assumes $f^{\theta}\in L^{2}(0,T;L^{2}(\mathcal{D}))$ and $u_{0}\in H^{1}(\mathcal{D})$.
@@
The triple $(u^{\varepsilon},\theta^{\varepsilon},\tau^{\varepsilon})$ is a solution of (1.1) on $[0,T]$ if $\tau^{\varepsilon}\colon\Omega\to[0,T]$ is a stopping time and $(u^{\varepsilon},\theta^{\varepsilon})$ is a stochastic process with trajectories $\mathbb{P}$-a.s. in
@@
Let $T>0$, $s\in(0,\tfrac{1}{2})$ and $p>4$, and choose $\gamma,\lambda$ as in (2.8) and (2.11). Let $\overline{\delta}_{\theta}$ and $\overline{\delta}_{u}$ be given by (2.12). Let parameters $\delta_{\theta},\delta_{u}$ satisfy
$0<\delta_{\theta}<\overline{\delta}_{\theta},\qquad 0<\delta_{u}<\overline{\delta}_{u},$
@@
(2.15) $\delta_{u}\geq\max\left\{\delta_{\theta},\frac{1}{2}-s\right\},\qquad\delta_{u}<1-\frac{2}{p},$
which define a non-empty set for the given range of $s$ and $p$. Set
$\alpha_{\theta}:=\frac{1}{4}+\frac{\delta_{\theta}}{2},\qquad\beta_{\theta}:=\frac{1}{4}-\frac{\delta_{\theta}}{4},$
@@ 3.1. Stochastic linear problem with non-homogeneous Dirichlet boundary conditions
We start this section by proving the first part of Proposition 2.1, in particular the well-posedness and the regularity for the mild solution $Z_{t}^{\varepsilon}$ of the linear problem (2.4) subject to Dirichlet noise boundary conditions. Note that this result is well-known in the literature, see for instance [10, Proposition 3.1].
@@
then the thesis follows from the Kolmogorov test for Gaussian processes, see [12, Proposition 3.16]. Consider
@@
For the estimate of the second integral $I_{2}$, fix $0\leq t_{1}\leq t_{2}$. By the semigroup property,
$e^{(t_{2}-r)\Delta}-e^{(t_{1}-r)\Delta}=e^{(t_{1}-r)\Delta}(e^{(t_{2}-t_{1})\Delta}-I),\qquad 0\leq r\leq t_{1}.$
Let $\gamma\in(0,1]$ (to be chosen later). Using the commutation of $(-\Delta)^{\sigma}$ with $e^{t\Delta}$ and inserting $(-\Delta)^{\gamma}(-\Delta)^{-\gamma}$, we obtain
@@
(3.5) $\begin{split}\norm{(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}(e^{(t_{2}-r)\Delta}-e^{(t_{1}-r)\Delta})}_{\mathcal{L}(L^{2})}&=\|(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}e^{(t_{1}-r)\Delta}(e^{(t_{2}-t_{1})\Delta}-I)\|_{\mathcal{L}(L^{2})}\\ &=\|(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}+\gamma}e^{(t_{1}-r)\Delta}(-\Delta)^{-\gamma}(e^{(t_{2}-t_{1})\Delta}-I)\|_{\mathcal{L}(L^{2})}\\ &\leq\|(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}+\gamma}e^{(t_{1}-r)\Delta}\|\,\|(-\Delta)^{-\gamma}(e^{(t_{2}-t_{1})\Delta}-I)\|_{\mathcal{L}(L^{2})}.\end{split}$
For the first factor we use [25, Theorem 6.13(c)] and we get
(3.6) $\|(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}+\gamma}e^{(t_{1}-r)\Delta}\|_{\mathcal{L}(L^{2})}\leq C(t_{1}-r)^{-(1-\alpha_{\theta}-\beta_{\theta}+\gamma)},$
for a constant $C$ independent of $t_{1}$. For the second factor, let $g\in L^{2}(\mathcal{D})$ and set $x:=(-\Delta)^{-\gamma}g\in D((-\Delta)^{\gamma})$. Since $(-\Delta)^{-\gamma}$ commutes with the heat semigroup $e^{(t_{2}-t_{1})\Delta}$, we have
$(-\Delta)^{-\gamma}(e^{(t_{2}-t_{1})\Delta}-I)g=(e^{(t_{2}-t_{1})\Delta}-I)(-\Delta)^{-\gamma}g=(e^{(t_{2}-t_{1})\Delta}-I)x.$
Applying [25, Theorem 6.13(d)] with exponent $\gamma$ yields
$\|(e^{(t_{2}-t_{1})\Delta}-I)x\|_{2}\leq C(t_{2}-t_{1})^{\gamma}\|(-\Delta)^{\gamma}x\|_{2}=C(t_{2}-t_{1})^{\gamma}\|g\|_{2},$
for a new constant $C$ independent of $t_{1},t_{2}$. Therefore
(3.7) $\|(-\Delta)^{-\gamma}(e^{(t_{2}-t_{1})\Delta}-I)\|_{\mathcal{L}(L^{2})}\leq C(t_{2}-t_{1})^{\gamma}.$
Combining (3.5)-(3.7) we obtain, for all $0\leq r\leq t_{1}$,
(3.8) $\|(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}(e^{(t_{2}-r)\Delta}-e^{(t_{1}-r)\Delta})\|_{\mathcal{L}(L^{2})}\leq C(t_{2}-t_{1})^{\gamma}(t_{1}-r)^{-(1-\alpha_{\theta}-\beta_{\theta}+\gamma)}.$
Using independence of $(\beta_{k})_{k}$, Itô isometry, and (3.8), we get
$\mathbb{E}\|I_{2}\|_{2}^{2}$ $=\sum_{k}\lambda_{k}^{2}\int_{0}^{t_{1}}\|(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}(e^{(t_{2}-r)\Delta}-e^{(t_{1}-r)\Delta})(-\Delta)^{\beta_{\theta}}De_{k}\|_{2}^{2}\,dr$
$\leq\sum_{k}\lambda_{k}^{2}\int_{0}^{t_{1}}\|(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}(e^{(t_{2}-r)\Delta}-e^{(t_{1}-r)\Delta})\|_{\mathcal{L}(L^{2})}^{2}\|(-\Delta)^{\beta_{\theta}}De_{k}\|_{2}^{2}\,dr$
@@
$\leq C(t_{2}-t_{1})^{2\gamma}\left(\sum_{k}\lambda_{k}^{2}\|(-\Delta)^{\beta_{\theta}}De_{k}\|_{2}^{2}\right)\int_{0}^{t_{1}}(t_{1}-r)^{-2(1-\alpha_{\theta}-\beta_{\theta}+\gamma)}\,dr.$
The time-integral is finite provided $2(1-\alpha_{\theta}-\beta_{\theta}+\gamma)<1$, i.e.
$\gamma<\alpha_{\theta}+\beta_{\theta}-\frac{1}{2}.$
Since $\alpha_{\theta}+\beta_{\theta}=\frac{1}{2}+\frac{\delta_{\theta}}{4}$, we may choose any $\gamma\in(0,\tfrac{\delta_{\theta}}{4})$, and then
(3.9) $\mathbb{E}\|I_{2}\|_{2}^{2}\leq C\,(t_{2}-t_{1})^{2\gamma}\sum_{k}\lambda_{k}^{2}\|(-\Delta)^{\beta_{\theta}}De_{k}\|_{2}^{2},$
@@
By the factorisation method, see [12, Section 5.3.1], since $\int_{0}^{T}t^{-2\rho}\|e^{t\Delta}\Phi\|_{L^{0}_{2}}^{2}dt<\infty$, it holds